OpenStudy (anonymous):

What is the value of the expression? -4 + (-6) times (-3) A. -22 B. -14 C.14 D.30

OpenStudy (anonymous):

@tester97 @RosieF

OpenStudy (anonymous):

14

OpenStudy (anonymous):

BEDMAS BRAH

OpenStudy (anonymous):

-4-6*-3 -6*-3 = 18 -4+18 = 14

OpenStudy (anonymous):

I got 30

OpenStudy (anonymous):

Believe what you like.

OpenStudy (tester97):

remember PEMDAS Parentheses Exponents Multiplication Division Addition Subtraction So therefore \[-4 + (-6)\times(-3)\] \[-4-6\times-3\] \[-4+18\] \[14\] So your answer here is C :)

OpenStudy (tester97):

Thats what i said.......

OpenStudy (tester97):

it took me forever but i still said it.......

OpenStudy (anonymous):

answer is C